If the class average is 195 and the SD is 30, what would be the score 1 SD above and below the mean?
Answer by jim_thompson5910(35256)

Score 1 SD Above the mean = mean + 1*SD = 195+1*30 = 225

Score 1 SD Below the mean = mean - 1*SD = 195-1*30 = 165